# Multiscale scattered data analysis in samplet coordinates

### v1

@@ 3. Condition number bound for generalized Vandermonde matrices
For every $\ell\in\mathbb{N}$ the condition number $\operatorname{cond}_{2}(\boldsymbol{K}_{\ell,\ell})$ is bounded independently of $N_{\ell}$ and $\delta_{\ell}$.
@@ 4. Samplets for scattered data approximation
Samplets are discrete signed measures, which exhibit vanishing moments. We briefly recall their underlying concepts as introduced in [7]. To this end, we consider the image of the spaces $\mathcal{S}_{\ell}$, $\ell=1,\ldots,L$, under the Riesz isomorphism $\mathcal{J}\colon\mathcal{N}_{\Phi_{\ell}}\to\mathcal{N}^{\prime}_{\Phi_{\ell}}$ and define
@@
by the identity $\mathcal{J}\Phi_{\ell}(\cdot-{\boldsymbol{x}})=\delta_{\boldsymbol{x}}$, where $\delta_{\boldsymbol{x}}$ is the point evaluation functional at ${\boldsymbol{x}}\in\Omega$. On the spaces $\mathcal{S}_{\ell}^{\prime}$, we introduce an inner product different from the canonical one by the defining property $\langle\delta_{{\boldsymbol{x}}_{i}},\delta_{{\boldsymbol{x}}_{j}}\rangle_{\mathcal{S}_{\ell}^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\delta_{ij}$ for ${\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j}\in X_{\ell}$.
For the sake of a more lightweight notation, we drop the subscript $\ell$ of the spaces $\mathcal{S}_{\ell}^{\prime}$ for the remainder of this paragraph and remark that each space $\mathcal{S}_{\ell}^{\prime}$ will in general have a different multiresolution analysis, where the maximum level $J$ depends on $\ell$. Given a multiresolution analysis
(14) $\mathcal{V}_{0}^{\prime}\subset\mathcal{V}_{1}^{\prime}\subset\cdots\subset\mathcal{V}_{J}^{\prime}=\mathcal{S}^{\prime},$
we keep track of the increment of information between two consecutive levels $j$ and $j+1$. Since there holds $\mathcal{V}_{j}^{\prime}\subset\mathcal{V}_{j+1}^{\prime}$, we may orthogonally decompose
$\mathcal{V}_{j+1}^{\prime}=\mathcal{V}_{j}^{\prime}\overset{\perp}{\oplus}\mathcal{W}_{j}^{\prime}$
for a certain detail space $\mathcal{W}_{j}^{\prime}$. In analogy to wavelet nomenclature, we call the elements of a basis of $\mathcal{V}_{0}^{\prime}$ scaling distributions and the elements of a basis of one of the spaces $\mathcal{W}_{j}^{\prime}$ samplets. This name is motivated by the idea that the basis distributions in $\mathcal{W}_{j}^{\prime}$ are supported at a small subsample or samplet of the data sites in $X$. The collection of the bases of $\mathcal{W}_{j}^{\prime}$ for $j=0,\ldots,J-1$ together with a basis of $\mathcal{V}_{0}^{\prime}$ is called a samplet basis for $\mathcal{S}^{\prime}$. A samplet basis can be constructed such that it exhibits vanishing moments of order $q+1$, i.e.,
(15) $(p,\sigma_{j,k})_{\Omega}=0\ \text{for all}\ p\in\mathcal{P}_{q},$
@@
The samplet construction for $\mathcal{S}^{\prime}$ is based on a cluster tree for the set of data sites $X$, i.e., a tree $\mathcal{C}$ with root $X$ such that each node $\tau\in\mathcal{C}$ is the disjoint union of its children. Such a cluster tree $\mathcal{C}$ directly induces a support based hierarchical clustering of the subspace $\mathcal{S}^{\prime}$, spanned by the Dirac-$\delta$-distributions supported at the data sites in $X$. With a slight abuse of notation, we will refer to this cluster tree also by $\mathcal{C}$ and to its elements by $\tau$. Given a cluster tree $\mathcal{C}$, a samplet basis for $\mathcal{S}^{\prime}$, where $\dim\mathcal{S}^{\prime}=N$, can be constructed with cost $\mathcal{O}(N)$. Assuming furthermore, that $\mathcal{C}$ is a balanced binary tree with maximum level $J=\lceil\log_{2}(N)\rceil$, the samplet basis exhibits the following properties, see [7] but also [6, 17].
@@
    The coefficient vector ${\boldsymbol{\omega}}_{j,k}=\big{[}\omega_{j,k,i}\big{]}_{i}$ of the samplet $\sigma_{j,k}$ satisfies the bound
    $\|{\boldsymbol{\omega}}_{j,k}\|_{1}\leq c2^{(J-j)/2}$
    for a constant $20<c<$ 2.
@@
A direct consequence of the orthonormality of the samplet basis is that the samplet transform
$[\sigma_{j,k}]_{j,k}={\boldsymbol{T}}[\delta_{{\boldsymbol{x}}_{i}}]_{i=1}^{N}$
satisfies ${\boldsymbol{T}}^{\intercal}{\boldsymbol{T}}={\boldsymbol{T}}{\boldsymbol{T}}^{\intercal}={\boldsymbol{I}}\in\mathbb{R}^{N\times N}$. If the samplet transform ${\boldsymbol{T}}$ is computed recursively adhering the hierarchy induced by the cluster tree $\mathcal{C}$, its cost is of order $\mathcal{O}(N)$.
Employing the Riesz isometry, the samplet basis gives rise to a basis for $\mathcal{S}\subset\mathcal{N}_{\Phi}$. To this end, consider a samplet
$\sigma_{j,k}=\sum_{i=1}^{N}\omega_{j,k,i}\delta_{{\boldsymbol{x}}_{i}}.$
The samplet can be identified with the function
$\psi_{j,k}\mathrel{\mathrel{\mathop{:}}=}\sum_{i=1}^{N}\omega_{j,k,i}\Phi({\boldsymbol{x}}_{i}-\cdot)\in\mathcal{S}$
by means of the Riesz isometry. The vanishing moment property (15) then translates to
$\langle\psi_{j,k},f\rangle_{\mathcal{N}_{\Phi}}=0$
for any $f\in\mathcal{N}_{\Phi}$ which satisfies $f|_{O}\in\mathcal{P}_{q}$ for any open and convex set $O$ with $\operatorname{supp}(\sigma_{j,k})\subset O\cap\Omega$.
@@
This means that the Gramian of the embedded samplet basis in $\mathcal{S}$ is just the samplet transformed generalized Vandermonde matrix. For asymptotically smooth kernels $K$, i.e. that there exist $C,r>0$ such that for all $({\boldsymbol{x}},{\boldsymbol{y}})\in(\Omega\times\Omega)\setminus\Delta$

### v2

@@ 3. Condition number bound for generalized Vandermonde matrices
For every $\ell\in\mathbb{N}$ the condition number $\operatorname{cond}_{2}(\boldsymbol{K}_{\ell,\ell})$ is bounded independently of $N_{\ell}$ and $\delta_{\ell}$. Furthermore, introducing the block diagonal scaling $\mathcal{D}\mathrel{\mathrel{\mathop{:}}=}\operatorname{diag}(\delta_{1}^{d}{\boldsymbol{I}},\ldots,\delta_{L}^{d}{\boldsymbol{I}}),$ the diagonally scaled matrix $\mathcal{D}^{1/2}\mathcal{K}\mathcal{D}^{1/2}$, cp. (10) is well conditioned, i.e., the corresponding condition number satisfies $\operatorname{cond}_{2}(\mathcal{D}^{1/2}\mathcal{K}\mathcal{D}^{1/2})\sim\operatorname{cond}_{F}(\mathcal{D}^{1/2}\mathcal{K}\mathcal{D}^{1/2})\sim 1$.
4. Numerical approximation of the multiscale system
In this paragraph, we discuss the approximation error issuing from replacing the matrix $\mathcal{K}$ of the multiscale system in (10) by an approximation $\widehat{\mathcal{K}}$.
As has been observed in Corollary 3.2, the multiscale system becomes well-conditioned by suitable diagonal scaling. For our analysis, we therefore consider the preconditioned matrix $\mathcal{K}_{\mathrm{pc}}\mathrel{\mathrel{\mathop{:}}=}\mathcal{D}^{1/2}\mathcal{K}\mathcal{D}^{1/2}$ and its numerical approximation $\widehat{\mathcal{K}}_{\mathrm{pc}}\mathrel{\mathrel{\mathop{:}}=}\mathcal{D}^{1/2}\widehat{\mathcal{K}}\mathcal{D}^{1/2}$. We have the following standard perturbation result, see, e.g., [12].
Lemma 4.1.
Let $\widehat{\mathcal{K}}$ be sufficiently close to $\mathcal{K}$ and let $\mathcal{K}_{\mathrm{pc}}{\boldsymbol{d}}=\mathcal{D}^{1/2}{\boldsymbol{f}}$ and $\widehat{\mathcal{K}}_{\mathrm{pc}}\widehat{\boldsymbol{d}}=\mathcal{D}^{1/2}{\boldsymbol{f}}$. Then there exists a constant $C>0$ such that the relative consistency error satisfies
$\frac{\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}}{\|{\boldsymbol{d}}\|_{2}}\leq C\frac{\|\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\|_{F}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}}.$
Proof.
The proof follows similarly to the proof of [12, Theorem 7.2]. There holds
$\mathcal{K}_{\mathrm{pc}}(\widehat{\boldsymbol{d}}-{\boldsymbol{d}})=-(\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}){\boldsymbol{d}}+(\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}})({\boldsymbol{d}}-\widehat{\boldsymbol{d}})$
Hence, by multiplying by $\mathcal{K}_{\mathrm{pc}}^{-1}$, taking norms and exploiting the consistency of the Frobenius and the Euclidean norm, we obtain
$\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}\leq\|\mathcal{K}_{\mathrm{pc}}^{-1}\|_{F}\|\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}\|_{F}\|{\boldsymbol{d}}\|_{2}+\|\mathcal{K}_{\mathrm{pc}}^{-1}\|_{F}\|\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}\|_{F}\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}.$
Factoring out common terms and simplifying the expression then results in
$\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}\leq\frac{\|\mathcal{K}_{\mathrm{pc}}^{-1}\|_{F}\|\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}\|_{F}}{1-\|\mathcal{K}_{\mathrm{pc}}^{-1}\|_{F}\|\widehat{\mathcal{K}}_{\mathrm{pc}}-\mathcal{K}_{\mathrm{pc}}\|_{F}}\|{\boldsymbol{d}}\|_{2}.$
Dividing by $\|{\boldsymbol{d}}\|_{2}$ and expanding the respective terms in the numerator and denominator by $\|\mathcal{K}_{\mathrm{pc}}\|_{F}$ finally gives
$\frac{\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}}{\|{\boldsymbol{d}}\|_{2}}\leq\frac{\operatorname{cond}_{F}(\mathcal{K}_{\mathrm{pc}})}{1-\operatorname{cond}_{F}(\mathcal{K}_{\mathrm{pc}})\|\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\|_{F}/\|\mathcal{K}_{\mathrm{pc}}\|_{F}}\frac{\|\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\|_{F}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}}.$
For a sufficiently small consistency error, we arrive at the claimed bound using Corollary 3.2. ∎
To connect the previous error estimate to the consistency error of the multiscale interpolant, we require the following technical lemma.
Lemma 4.2.
Consider any function $v_{\ell}=\sum_{i=1}^{N_{\ell}}a_{i}\Phi_{\ell}(\cdot-{\boldsymbol{x}}_{i}^{(\ell)})\in\mathcal{S}_{\ell}$. Then, there holds
$\|v_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}\sim\delta_{\ell}^{-d/2}\|{\boldsymbol{a}}\|_{2}\quad\text{as well as}\quad\|v_{\ell}|X_{\ell}\|_{\ell^{2}}\sim\delta_{\ell}^{-d}\|{\boldsymbol{a}}\|_{2},$
where ${\boldsymbol{a}}\mathrel{\mathrel{\mathop{:}}=}[a_{1},\ldots,a_{N_{\ell}}]^{\intercal}$ is the coefficient vector of $v_{\ell}$.
Proof.
By the reproducing property and due to $\|{\boldsymbol{K}}_{\ell,\ell}\|_{2}\sim\|{\boldsymbol{K}}_{\ell,\ell}^{-1}\|_{2}\sim\delta^{-d}_{\ell}$, we obtain
$\|v_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}={\boldsymbol{v}}^{\intercal}{\boldsymbol{K}}_{\ell,\ell}{\boldsymbol{a}}\sim\|{\boldsymbol{K}}_{\ell,\ell}\|_{2}\|{\boldsymbol{a}}\|_{2}^{2}\sim\delta^{-d}_{\ell}\|{\boldsymbol{a}}\|_{2}^{2}.$
Taking square roots yields the first claim. Similarly, we observe
$\|v_{\ell}|X_{\ell}\|_{\ell^{2}}^{2}=\|{\boldsymbol{K}}_{\ell,\ell}{\boldsymbol{a}}\|_{2}^{2}={\boldsymbol{a}}^{\intercal}{\boldsymbol{K}}_{\ell,\ell}^{2}{\boldsymbol{a}}\sim\delta_{\ell}^{-2d}\|{\boldsymbol{a}}\|_{2}^{2}$
and the second claim is again obtained by taking square roots. ∎
Combining the previous two lemmata, we obtain a bound on the consistency error of the residual corrections.
Theorem 4.3.
Let $w_{\ell}=\sum_{i=1}^{N_{\ell}}c_{\ell,i}\Phi_{\ell}(\cdot-{\boldsymbol{x}}_{i})\in\mathcal{S}_{\ell}$ for $\ell=1,\ldots,L$ denote the residual corrections obtained from the solution of $\mathcal{K}{\boldsymbol{c}}={\boldsymbol{f}}$, see (10), and let $\widehat{w}_{\ell}=\sum_{i=1}^{N_{\ell}}\widehat{c}_{\ell,i}\Phi_{\ell}(\cdot-{\boldsymbol{x}}_{i})\in\mathcal{S}_{\ell}$ for $\ell=1,\ldots,L$ be the residual corrections obtained from solving the perturbed system $\widehat{\mathcal{K}}\widehat{\boldsymbol{c}}={\boldsymbol{f}}$. Then, there holds
$\frac{\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}}{\sum_{\ell=1}^{L}\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}}\leq C\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}^{2}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}^{2}}$
for a constant $C>0$.
Proof.
Since the solutions to the original systems and their preconditioned versions are connected via ${\boldsymbol{d}}=\mathcal{D}^{-1/2}{\boldsymbol{c}}$ and $\widehat{\boldsymbol{d}}=\mathcal{D}^{-1/2}\widehat{\boldsymbol{c}}$, respectively, the first equivalence in Lemma 4.2 yields
$\|w_{\ell}-\widehat{w}_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}\sim\delta^{-d}_{\ell}\|{\boldsymbol{c}}_{\ell}-\widehat{\boldsymbol{c}}_{\ell}\|_{2}^{2}=\|{\boldsymbol{d}}_{\ell}-\widehat{\boldsymbol{d}}_{\ell}\|_{2}^{2}$
as well as
$\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}\sim\delta^{-d}_{\ell}\|{\boldsymbol{c}}_{\ell}\|_{2}^{2}=\|{\boldsymbol{d}}_{\ell}\|_{2}^{2}.$
In view of $\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}^{2}=\sum_{\ell=1}^{L}\|{\boldsymbol{d}}_{\ell}-\widehat{\boldsymbol{d}}_{\ell}\|_{2}^{2}$ and $\|{\boldsymbol{d}}\|_{2}^{2}=\sum_{\ell=1}^{L}\|{\boldsymbol{d}}_{\ell}\|_{2}^{2}$, we arrive at
$\frac{\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}}{\sum_{\ell=1}^{L}\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}}\sim\frac{\|{\boldsymbol{d}}-\widehat{\boldsymbol{d}}\|_{2}^{2}}{\|{\boldsymbol{d}}\|_{2}^{2}}\leq C\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}^{2}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}^{2}}$
by employing Lemma 4.1. This completes the proof. ∎
To bound the numerical approximation error between $s_{L}$ and $\widehat{s}_{L}\mathrel{\mathrel{\mathop{:}}=}\widehat{w}_{1}+\dots+\widehat{w}_{L}$, we have the following result, which links this error to the approximation error of the residual corrections.
Lemma 4.4.
Under the conditions of Theorem 4.3, the approximation $\widehat{s}_{L}\mathrel{\mathrel{\mathop{:}}=}\widehat{w}_{1}+\dots+\widehat{w}_{L}$ to $s_{L}$ satisfies the estimate
$\|s_{L}-\widehat{s}_{L}\|_{L^{2}(\Omega)}^{2}\leq CL\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}.$
for some constant $C>0$.
Proof.
From the Cauchy–Schwarz inequality, it follows that
$\|s_{L}-\widehat{s}_{L}\|^{2}_{L^{2}(\Omega)}=\bigg{\|}\sum_{\ell=1}^{L}(w_{\ell}-\widehat{w}_{\ell})\bigg{\|}^{2}_{L^{2}(\Omega)}\leq L\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{L^{2}(\Omega)}.$
Next, we consider the sampling inequality from [16, Theorem 3.5], i.e.,
$\|u\|_{L^{p}}\leq C\Big{(}h_{X,\Omega}^{\theta}\|u\|_{H^{\theta}(\Omega)}+h_{X,\Omega}^{d/p}\|u|X\|_{\ell^{2}}\Big{)},\quad C>0.$
Letting $p=2$ and squaring the previous inequality results in
$\|u\|_{L^{2}(\Omega)}^{2}\leq 2C\Big{(}h_{X,\Omega}^{2\theta}\|u\|_{H^{\theta}(\Omega)}^{2}+h_{X,\Omega}^{d}\|u|X\|_{\ell^{2}}^{2}\Big{)}.$
Inserting this in the first inequality, we arrive at
$\|s_{L}-\widehat{s}_{L}\|^{2}_{L^{2}(\Omega)}\leq 2CL\sum_{\ell=1}^{L}\big{(}h_{\ell}^{2\theta}\|w_{\ell}-\widehat{w}_{\ell}\|_{H^{\theta}(\Omega)}^{2}+h_{\ell}^{d}\|w_{\ell}-\widehat{w}_{\ell}|X_{\ell}\|_{\ell^{2}}^{2}\big{)}.$
Next, we have from [29, Lemma 1] that
$\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{H^{\theta}(\Omega)}\leq C\delta_{\ell}^{-2\theta}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}},\quad C>0,$
while the second equivalence in Lemma 4.2 amounts to
$\|w_{\ell}-\widehat{w}_{\ell}|X_{\ell}\|_{\ell^{2}}^{2}\sim\delta_{\ell}^{-2d}\|{\boldsymbol{c}}_{\ell}-\widehat{\boldsymbol{c}}_{\ell}\|_{2}^{2}\sim\delta_{\ell}^{-d}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}.$
Inserting these two estimates in the previous one yields with $\delta_{\ell}=\eta h_{\ell}$, cp. (4), that
$\|s_{L}-\widehat{s}_{L}\|^{2}_{L^{2}(\Omega)}\leq CL\sum_{\ell=1}^{L}\big{(}\eta^{-2\theta}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}+\eta^{-d}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}\big{)},$
for some constant $C>0$, from which the claim is immediately obtained. ∎
Finally, we need a bound on the residual corrections by the right hand side $f$.
Lemma 4.5.
Under the conditions of Theorem 2.1, there exists a constant $C>0$ such that
$\sum_{\ell=1}^{L}\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}\leq C\|f\|^{2}_{H^{\theta}(\Omega)}.$
Proof.
By construction, the residual correction $w_{\ell}$ is the norm-minimal interpolant to the residual $e_{\ell-1}\mathrel{\mathrel{\mathop{:}}=}f-\sum_{m=1}^{\ell-1}w_{m}$, for any $1\leq\ell\leq L$, which implies
$\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}\leq\|Ee_{\ell-1}||_{\mathcal{N}_{\Phi_{\ell}}}$
with the continuous extension operator $E\colon H^{\theta}(\Omega)\to H^{\theta}(\mathbb{R}^{d})$, $\|Ef\|_{H^{\theta}(\mathbb{R}^{d})}\leq C_{E}\|f\|_{H^{\theta}(\Omega)}$. Now we can use the recursion estimate [29, Theorem 1] to obtain
$\|Ee_{\ell-1}||_{\mathcal{N}_{\Phi_{\ell}}}\leq(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})\|Ee_{\ell-2}\|_{\mathcal{N}_{\Phi_{\ell-1}}}.$
Recurring $\ell-2$ times, we arrive at the bound
$\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}\leq\|Ee_{\ell-1}||_{\mathcal{N}_{\Phi_{\ell}}}$ $\leq(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{\ell-2}\|Ee_{0}\|_{\mathcal{N}_{\Phi_{1}}}$
$=(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{\ell-2}\|Ef\|_{\mathcal{N}_{\Phi_{1}}}$
$\leq C_{E}(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{\ell-2}\delta_{1}^{-\theta}\|f\|_{H^{\theta}(\Omega)}.$
By this bound and by employing the summation formula for the geometric series, we finally estimate
$\sum_{\ell=1}^{L}\|w_{\ell}\|_{\mathcal{N}_{\Phi_{\ell}}}^{2}$ $\leq\sum_{\ell=1}^{L}C_{E}^{2}c(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2(\ell-2)}\delta_{1}^{-2\theta}\|f\|_{H^{\theta}(\Omega)}^{2}$
$=\frac{C_{E}^{2}\delta_{1}^{-2\theta}}{(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2}}\|f\|_{H^{\theta}(\Omega)}^{2}\sum_{\ell=0}^{L-1}(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2\ell}$
$=\frac{C_{E}^{2}\delta_{1}^{-2\theta}}{(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2}}\frac{1-(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2L}}{1-(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{2}}\|f\|_{H^{\theta}(\Omega)}^{2}.$
This completes the proof. ∎
Combining the previous lemmata with Theorem 2.1, we arrive at the final error estimate.
Theorem 4.6.
Under the conditions of Theorems 2.1 and 4.3, there exists a constant $C>0$ such that
$\|f-\widehat{s}_{L}\|_{L^{2}(\Omega)}\leq C\Bigg{(}(C_{1}\mu^{\theta}+C_{1}\eta^{-\theta})^{L}+\sqrt{L}\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}}\Bigg{)}\|f\|_{H^{\theta}(\Omega)}.$
Proof.
By the triangle inequality, there holds
$\|f-\widehat{s}_{L}\|_{L^{2}(\Omega)}\leq\|f-{s}_{L}\|_{L^{2}(\Omega)}+\|s_{L}-\widehat{s}_{L}\|_{L^{2}(\Omega)}.$
The first term on the right hand side is bounded using Theorem 2.1. For the second term, we obtain using Lemmata 4.4 and 4.5 as well as Theorem 4.3 that
$\|s_{L}-\widehat{s}_{L}\|_{L^{2}(\Omega)}^{2}$ $\leq CL\sum_{\ell=1}^{L}\|w_{\ell}-\widehat{w}_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}\leq CL\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}^{2}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}^{2}}\sum_{\ell=1}^{L}\|w_{\ell}\|^{2}_{\mathcal{N}_{\Phi_{\ell}}}$
$\leq CL\frac{\big{\|}\mathcal{K}_{\mathrm{pc}}-\widehat{\mathcal{K}}_{\mathrm{pc}}\big{\|}_{F}^{2}}{\|\mathcal{K}_{\mathrm{pc}}\|_{F}^{2}}\|f\|_{H^{\theta}(\Omega)}^{2}.$
Taking square roots and inserting into the first inequality yields the claim. ∎
@@ 5. Samplets for scattered data approximation
Samplets are discrete signed measures, which exhibit vanishing moments. We briefly recall their underlying concepts as introduced in [9]. To this end, we consider the image of the spaces $\mathcal{S}_{\ell}$, $\ell=1,\ldots,L$, defined in Equation (7), under the Riesz isomorphism $\mathcal{J}\colon\mathcal{N}_{\Phi_{\ell}}\to\mathcal{N}^{\prime}_{\Phi_{\ell}}$ and define
@@
by the identity $\mathcal{J}\Phi_{\ell}(\cdot-{\boldsymbol{x}})=\delta_{\boldsymbol{x}}$, where $\delta_{\boldsymbol{x}}$ is the point evaluation functional at ${\boldsymbol{x}}\in\Omega$. We equip the spaces $\mathcal{S}_{\ell}^{\prime}$ with an inner product, different from the canonical one, by the defining property $\langle\delta_{{\boldsymbol{x}}_{i}},\delta_{{\boldsymbol{x}}_{j}}\rangle_{\mathcal{S}_{\ell}^{\prime}}\mathrel{\mathrel{\mathop{:}}=}\delta_{ij}$ for ${\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j}\in X_{\ell}$.
For the sake of a more lightweight notation, we drop the subscript $\ell$ of the spaces $\mathcal{S}_{\ell}^{\prime}$ for the remainder of this paragraph and remark that each space $\mathcal{S}_{\ell}^{\prime}$ will in general have a different multiresolution analysis, where the maximum level $J$ depends on $\ell$. Given a multiresolution analysis $\mathcal{V}_{0}^{\prime}\subset\mathcal{V}_{1}^{\prime}\subset\cdots\subset\mathcal{V}_{J}^{\prime}=\mathcal{S}^{\prime},$ we keep track of the increment of information between two consecutive levels $j$ and $j+1$. Since there holds $\mathcal{V}_{j}^{\prime}\subset\mathcal{V}_{j+1}^{\prime}$, we may orthogonally decompose $\mathcal{V}_{j+1}^{\prime}=\mathcal{V}_{j}^{\prime}{\oplus}\mathcal{W}_{j}^{\prime}$ for a certain detail space $\mathcal{W}_{j}^{\prime}\perp\mathcal{V}_{j}^{\prime}$. In analogy to wavelet nomenclature, we call the elements of a basis of $\mathcal{V}_{0}^{\prime}$ scaling distributions and the elements of a basis of one of the spaces $\mathcal{W}_{j}^{\prime}$ samplets. This name is motivated by the idea that the basis distributions in $\mathcal{W}_{j}^{\prime}$ are supported at a small subsample or samplet of the data sites in $X$. The collection of the bases of $\mathcal{W}_{j}^{\prime}$ for $j=0,\ldots,J-1$ together with a basis of $\mathcal{V}_{0}^{\prime}$ is called a samplet basis for $\mathcal{S}^{\prime}$. A samplet basis can be constructed such that it exhibits vanishing moments of order $q+1$, i.e.,
(15) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\sigma_{j,k}(p)=0\quad\text{for all}\ p\in\mathcal{P}_{q},}$
@@
The samplet construction for $\mathcal{S}^{\prime}$ is based on a cluster tree for the set of data sites $X$, i.e., a tree $\mathcal{C}$ with root $X$ such that each node $\tau\in\mathcal{C}$ is the disjoint union of its children. Such a cluster tree $\mathcal{C}$ directly induces a support based hierarchical clustering of the subspace $\mathcal{S}^{\prime}$, spanned by the Dirac-$\delta$-distributions supported at the data sites in $X$. With a slight abuse of notation, we will refer to this cluster tree also by $\mathcal{C}$ and to its nodes by $\tau$. Given a cluster tree $\mathcal{C}$, a samplet basis for $\mathcal{S}^{\prime}$, where $\dim\mathcal{S}^{\prime}=N$, can be constructed with cost $\mathcal{O}(N)$. Assuming furthermore, that $\mathcal{C}$ is a balanced binary tree with maximum level $J=\lceil\log_{2}(N)\rceil$, the samplet basis exhibits the following properties, see [9] but also [8, 25].
@@
    The coefficient vector ${\boldsymbol{\omega}}_{j,k}=\big{[}\omega_{j,k,i}\big{]}_{i}$ of the samplet $\sigma_{j,k}$ satisfies the bound $\|{\boldsymbol{\omega}}_{j,k}\|_{1}\leq c2^{(J-j)/2}$ for a constant $20<c<$ 2.
@@
A direct consequence of the orthonormality of the samplet basis is that the samplet transform $[\sigma_{j,k}]_{j,k}={\boldsymbol{T}}[\delta_{{\boldsymbol{x}}_{i}}]_{i=1}^{N}$ satisfies ${\boldsymbol{T}}^{\intercal}{\boldsymbol{T}}={\boldsymbol{T}}{\boldsymbol{T}}^{\intercal}={\boldsymbol{I}}\in\mathbb{R}^{N\times N}$. If the samplet transform ${\boldsymbol{T}}$ is computed recursively adhering the hierarchy induced by the cluster tree $\mathcal{C}$, its cost is of order $\mathcal{O}(N)$.
Employing the Riesz isometry, the samplet basis induces a basis for $\mathcal{S}\subset\mathcal{N}_{\Phi}$. To this end, consider a samplet $\sigma_{j,k}=\sum_{i=1}^{N}\omega_{j,k,i}\delta_{{\boldsymbol{x}}_{i}}.$ The samplet can be identified with the function $\psi_{j,k}\mathrel{\mathrel{\mathop{:}}=}\sum_{i=1}^{N}\omega_{j,k,i}\Phi({\boldsymbol{x}}_{i}-\cdot)\in\mathcal{S}$ by means of the Riesz isometry. The vanishing moment property (15) then translates to $\langle\psi_{j,k},f\rangle_{\mathcal{N}_{\Phi}}=0$ for any $f\in\mathcal{N}_{\Phi}$ which satisfies $f|_{\operatorname{supp}(\sigma_{j,k})}\in\mathcal{P}_{q}$.
@@
This means that the Gramian of the embedded samplet basis in $\mathcal{S}$ is just the samplet transformed generalized Vandermonde matrix. For asymptotically smooth kernels $K$, i.e., there exist $C,r>0$ such that for all $({\boldsymbol{x}},{\boldsymbol{y}})\in(\Omega\times\Omega)\setminus\Delta$